### v1

@@ 5.1.4 Hypersurface decision trees are proper
Proof.
Axioms 1 and 2 follow directly from the definition of $h^{\pm}$. Axioms 3 and 4 hold because, for any hyperplane $h$ defined by a $D$-tuple of data points $\mathit{xs}_{D}$ we have
@@
Thus axiom 3, 4 holds. ∎
@@
Proof.
We prove this by induction. The base case, involving a single linear hypersurface, has already been established in Theorem 12. For the inductive step, assume two subtrees $t_{1}$ and $t_{2}$ of size $i$ and Ki1K-i-1, respectively, are optimal for the corresponding subregions $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$, determined by root $h$. We show that the tree $t=DN\left(t_{1},h,t_{2}\right)$ is optimal with respect to root $h$. Since $t_{1}$ and $t_{2}$ are optimal, any subtree of the same size constructed in $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$ will have a equal or greater 0-1 loss than $t_{1}$ and $t_{2}$. By definition of the objective function, we have
@@
Thus, $t=DN\left(t_{1},h,t_{2}\right)$ is optimal for the fixed root $h$. Furthermore, for any arbitrary hypersurface defined by normal vector $\boldsymbol{w}\in\mathbb{R}^{G}$, there exists a hypersurface passing through $G$ points that induces the same partition $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$. Therefore, the optimal decision tree constructed using hypersurfaces through $G$ points attains the same 0-1 loss as the optimal tree constructed from arbitrary hypersurfaces. ∎
@@ 5.2.1 Generating fixed polynomial degree hypersurface splitting rules
    $C_{1}=\left(\begin{array}[]{c}n\\ G\end{array}\right)$, $C_{2}=\left(\begin{array}[]{c}n+1\\ G\end{array}\right)$
@@
To solve the original decision tree problem $\mathit{odt}_{\text{size}}$ using $\mathit{sodt}$, a separate generator is required to produce possible splitting rules for a given data list $\mathit{xs}$. Specifically, we need an efficient definition for the following specification:
$\mathit{odk}_{\text{size}}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}$
where $\mathit{gen}_{\mathit{splits}}^{M}:\mathcal{D}\to\left[\mathcal{R}\right]$ denotes the splitting rule generator, precisely the splitting rule generator used in the definition of $\mathit{odk}_{\text{depth}}$.
For the axis-parallel decision tree problem, $\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)$ can be simply defined as the all possible axis-parallel hyperplanes that go through every data points in $\mathit{xs}$, so there are $N\times D$ such hyperplanes in total. For more complex splitting rules, as established in Theorems 10 and 12, both hyperplane and hypersurface decision boundaries can be exhaustively enumerated by generating $G$-combinations of data points.
Consequently, the search space of hypersurface decision boundaries can be represented as $\mathcal{S}_{\text{kcombs}}\left(G,\mathit{ds}\right)$, and the generator for this search space can be implemented directly using a combination generator:
$\mathit{gen}_{\mathit{splits}}^{M}\left(\mathit{ds}\right)=!\left(G\right)\circ\mathit{kcombs}\left(G,\mathit{ds}\right),$
@@
$\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}=\mathit{kcombs}\left(K\right)\circ!\left(G\right)\circ\mathit{kcombs}\left(G,\mathit{ds}\right),$
@@
However, the naive implementation 47 is inefficient due to the overhead of storing intermediate results: $\mathit{kcombs}\left(G\right)$ produces $O\left(N^{G}\right)$ combinations for an input of size $N$, which is both memory-intensive and computationally costly. Additionally, its non-recursive nature prevents the application of thinning techniques.
Nevertheless, He (2025) presents both D&C and sequential versions of $\mathit{nestedCombs}$, while the D&C version is efficient, it complicates the ordering of elements. Element ordering is critical for our applications because combinations represent hyperplanes, which carry essential predictive information. Therefore, we adopt the sequential version of $\mathit{nestedCombs}$ developed by He and Little (2025a). The pseudo-code for this sequential version is presented in Algorithm 1.
@@ 5.2.2 Generating mixed polynomial degree hypersurface splitting rules
       $C_{1}=\left(\begin{array}[]{c}G\\ D\end{array}\right)$, $C_{2}=\left(\begin{array}[]{c}G+1\\ D\end{array}\right)$
@@
One advantage of constructing algorithms based on a generic framework is the flexibility it provides. In theory, since the input to $\mathit{odt}_{\text{size}}$ is a list of rules rather than raw data, arbitrary splitting rules satisfying the decision tree axioms can be supplied. This allows the construction of decision trees with mixed splitting rules. The main challenge lies in generating these splitting rules efficiently from a given data list. Fortunately, axis-parallel hyperplanes, general hyperplanes, and hypersurfaces can all be represented as combinations of data items.
A key observation about Algorithm 1 is that, it can generate not only $G$-combinations but all $G^{\prime}$-combinations for $0\leq G^{\prime}\leq G$. However, the nested combination generator $\mathit{nestedCombs}\left(K,G\right)$ uses only the $G$-combinations and ignores smaller $G^{\prime}$-combination ($G^{\prime}<G$). To accomodate this, we need to modify specification 47 to the following
@@
where $\left(\mathit{Gs},\mathit{xss}\right)=\mathit{concat}\left[!\left(G,\mathit{xss}\right)\mid G\leftarrow\mathit{Gs}\right]$ and $\mathit{Gs}=\left[G_{1},G_{2},\ldots G_{l}\right]$ is an ascending list of integers ($G_{1}<G_{2}\ldots<G_{l}$), with $G_{l}$ being the largest, corresponds to a list of polynomial degree $\mathit{Ms}=\left[M_{1},M_{2},\ldots M_{l}\right]$. Specification 48 selects several $G$-combinations, such that $G\in\mathit{Gs}$, and constructs $K$-nested combinations from them.
As usual, the naive specification 48 is inefficient. He and Little (2025a) provide an efficient solution by defining the recursive pattern of $\mathit{nestedCombs}^{\prime}\left(K,\mathit{Gs}\right)$ as
$\bigg{<}\mathit{setEmpty}\left(G_{l}\right)\cdot\mathit{KcombsAlg}\left(K\right)\cdot\mathit{Ffst},$
@@
$convol_{\text{new}}\left(\cup,\circ,\mathit{css}_{1},\mathit{css}_{2}\right)=\left[\mathit{concat}\left(\mathit{zipWith}\left(\circ,\mathit{init}\left(\mathit{tail}\left(\mathinner{css}\right)\right),\mathit{init}\left(\mathit{tail}\left(\mathit{css}_{2}\right)\right)\right)\right)\mid\mathinner{css}\leftarrow\mathit{inits}\left(\mathit{css}_{1}\right)\right].$ (49)
@@ 5.3.1 Complexity of size-constrained trees
Furthermore, the average-case complexity of $\mathit{sodt}_{\text{rec}}$ is substantially lower than its theoretical upper bound. The worst-case complexity of $\mathit{sodt}_{{}^{\prime}\text{rec}}$ occurs only when the tree forms a “chain” structure, which, as shown in the empirical analysis in Part II, is a rare occurrence.
@@ 5.3.2 Complexity of depth-constrained trees
$\mathcal{S}_{\text{depth}}\left(d,\mathit{xs}\right)\geq\mathcal{S}_{\text{size}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right).$ (53)
where $\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)$ denote the all possible unique splitting rules generated by $\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)$, as there might have duplicate splitting rules (for categorical data, distinct data points may have same feature values).
@@
$\mathcal{S}_{\text{size}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right)\geq\mathcal{S}_{\text{kcombs}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right).$ (54)
@@
$\mathcal{S}_{\text{depth}}\left(d,\mathit{xs}\right)\geq\mathcal{S}_{\text{size}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right)\geq\mathcal{S}_{\text{kcombs}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right)=\left(\begin{array}[]{c}\left|\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right|\\ d\end{array}\right)$ (55)
As discussed, for axis parallel decision tree problem, $\left|\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right|$ is the sum of all distinct features with respect to the training data list $\mathit{xs}$, which is at most $N\times D$ for a size $N$ data list in $\mathbb{R}^{D}$.
@@
In the original source code of Mazumder et al. (2022)’s Quant-BnB algorithm, there is a function count_remain_3D, which computes the combinatorial complexity of the search space, i.e., the number of remaining trees in each iteration. The initial call to count_remain_3D, calculates this complexity directly from the input data without bounding parameters. For almost all datasets we have tested, count_remain_3D computed results that is strictly smaller than $\left(\begin{array}[]{c}\left|\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right|\\ d\end{array}\right)$. Interested readers are encouraged to test this result on their own datasets.
As explained above, this should be impossible because every $d$-combination of splitting rule should construct at least one possible proper decision tree. We hypothesize that Mazumder et al. (2022) either employed an incorrect algorithmic procedure for solving the AODT problem or addressed a specialized ODT problem. However, their ambiguous problem definition hinders verification of their claims. Consequently, there is an urgent need for Mazumder et al. (2022) to clarify their counting formula to validate the correctness of their algorithms.

### v2

@@ 5.1.4 Hypersurface decision trees are proper
Proof Axioms 1 and 2 follow directly from the definition of $h^{\pm}$. Axioms 3 and 4 hold because, for any hyperplane $h$ defined by a $D$-tuple of data points $\mathit{xs}_{D}$ we have
@@
Thus axiom 3, 4 holds.
@@
Proof We prove this by induction. The base case, involving a single linear hypersurface, has already been established in Theorem 25. For the inductive step, assume two subtrees $t_{1}$ and $t_{2}$ of size $i$ and Ki1K-i-1, respectively, are optimal for the corresponding subregions $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$, determined by root $h$. We show that the tree $t=DN\left(t_{1},h,t_{2}\right)$ is optimal with respect to root $h$. Since $t_{1}$ and $t_{2}$ are optimal, any subtree of the same size constructed in $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$ will have a equal or greater 0-1 loss than $t_{1}$ and $t_{2}$. By definition of the objective function, we have
@@
Thus, $t=DN\left(t_{1},h,t_{2}\right)$ is optimal for the fixed root $h$. Furthermore, for any arbitrary hypersurface defined by normal vector $\boldsymbol{w}\in\mathbb{R}^{G}$, there exists a hypersurface passing through $G$ points that induces the same partition $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$. Therefore, the optimal decision tree constructed using hypersurfaces through $G$ points attains the same 0-1 loss as the optimal tree constructed from arbitrary hypersurfaces.
@@ 5.2 Splitting rule generators
Note that both $\mathit{odt}_{\text{size}}:\left[\mathcal{R}\right]\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)$ and $\mathit{odt}_{\text{depth}}:\mathcal{D}\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)$ cannot yet be directly applied to solve the optimal decision tree problem in machine learning. This is because we have not explained how to generate the input $\mathit{rs}:\left[\mathcal{R}\right]$ for $\mathit{odt}_{\text{size}}$ or the function $\mathit{gen}_{\mathit{splits}}:\mathcal{D}\to\left[\mathcal{R}\right]$ used in $\mathit{odt}_{\text{depth}}$. It turns out that both functions are the same function. In this section, we describe how to construct $\mathit{gen}_{\mathit{splits}}$ for solving various ML problems, including axis-parallel, hyperplane, hypersurface, and mixed-splitting ODT problems.
Specifically, the general paradigm for solving the size-constrained optimal decision tree problem requires composing $\mathit{odk}_{\text{size}}\left(K\right)$ with $\mathit{gen}_{\mathit{splits}}^{M}$. In other words, we obtain the following specification:
$\mathit{odt}_{\text{size}}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}$ (64)
where $\mathit{gen}_{\mathit{splits}}$ is precisely the splitting rule generator used in the definition of $\mathit{odk}_{\text{depth}}$.
In principle, as long as the rules $\mathit{rs}:\left[\mathcal{R}\right]$ generated by $\mathit{gen}_{\mathit{splits}}^{M}$ satisfy the decision tree axioms 1, $\mathit{odt}_{\text{size}}\left(K\right)$ can solve the size-constrained ODT problem exactly with respect to $\mathit{rs}$. This modularized design provides substantial flexibility for generalizing our algorithm to different problems. In practice, we need only redefine $\mathit{gen}_{\mathit{splits}}^{M}$ to adapt the program to new problems, while the main program $\mathit{odt}_{\text{size}}\left(K\right)$ remains unchanged, which demonstrates one of the primary advantage of adopting a general formalism in algorithm design.
For example, in the axis-parallel decision tree problem, $\mathit{gen}_{\mathit{splits}}^{0}:\mathcal{D}\to\left[\mathcal{H}^{0}\right]$ can be defined as the set of all possible axis-parallel hyperplanes passing through each data point in $\mathit{xs}:\mathcal{D}$. Thus, there are $N\times D$ such hyperplanes in total. More concretely, given a dataset $\mathit{xs}:\mathcal{D}$ in $\mathbb{R}^{D}$,
$\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)=\left[x_{i}\mid x_{i}\leftarrow x,x\leftarrow\mathit{xs}\right],$ (65)
By substituting $\mathit{gen}_{\mathit{splits}}^{0}$ into 1
$\mathit{aodt}_{K}=\mathit{odt}_{\text{size}}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{0},$
where $\mathit{aodt}$ is short for “axis-parallel optimal decision tree problem,” which yields an efficient solution for solving the classical optimal decision tree problem in machine learning.
Constructing generators for more complex splitting rules is more involved, but their structured form allows the composed function $\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}$ to be fused into a more efficient algorithm. In this section, we detail the construction of fused programs for both r $\mathit{gen}_{\mathit{splits}}^{M}$—the generator of hypersurface splitting rules—and $\mathit{gen}_{\mathit{splits}}^{m}$ ($0\leq m\leq M$) —the generator of mixed splitting rules that combine hypersurface, hyperplane, and axis-parallel rules.
@@ 5.2.1 Generating fixed polynomial degree hypersurface splitting rules
      $C_{1}=\left(\begin{array}[]{c}n\\ G-1\end{array}\right)$, $C_{2}=\left(\begin{array}[]{c}n\\ G\end{array}\right)$
@@
Due to the isomorphism between hyperplane classification and hypersurface classification established in Theorem 24, we can define the optimal hypersurface decision tree algorithm $\mathit{hodt}^{M}:\mathbb{N}\times\mathcal{D}\to\mathit{DTree}\left(\mathcal{H}^{M},\mathcal{D}\right)$ as the composition of $\mathit{odk}_{\text{size}}\left(K\right)$ with a hypersurface splitting rule generator $\mathit{gen}_{\mathit{splits}}^{M}$, together with an embedding function $\mathit{embed}_{M}$ (the programmatic definition of $\rho_{M}$). In other words, we obtain the following specification:
$\mathit{hodt}_{K}^{M}$ $=\mathit{odk}_{\text{size}}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}\circ\mathit{embed}_{M}$ (66)
$=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}\circ\mathit{embed}_{M}$
As established in Theorems 23 and 25, both hyperplane and hypersurface decision boundaries can be exhaustively enumerated by generating $G$-combinations of data points. Consequently, the search space of hypersurface decision boundaries can be represented as $\mathcal{S}_{\text{kcombs}}\left(G,\mathit{ds}\right)$, and the generator for this search space can be implemented directly using a combination generator:
$\mathit{gen}_{\mathit{splits}}^{M}\left(\mathit{ds}\right)=!\left(G\right)\circ\mathit{kcombs}\left(G,\mathit{ds}\right),$ (67)
@@
$\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}=\mathit{kcombs}\left(K\right)\circ!\left(G\right)\circ\mathit{kcombs}\left(G,\mathit{ds}\right),$ (68)
@@
Then we can rewrite 66 by following
$\mathit{hodt}_{K}^{M}$ $=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}\circ\mathit{embed}_{M}$
$\subseteq\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{nestedCombs}\left(K,G\right)\circ\mathit{embed}_{M}$
However, the naive implementation 69 is inefficient due to the overhead of storing intermediate results: $\mathit{kcombs}\left(G\right)$ produces $O\left(N^{G}\right)$ combinations for an input of size $N$, which is both memory-intensive and computationally costly. Additionally, its non-recursive nature prevents the application of thinning techniques.
Our goal is now to fuse the two-step definition of $\mathit{nestedCombs}\left(K,G\right)$, given in 69 into a single recursive function. He (2025b) presents both divide-and-conquer (D&C) and sequential versions for the fused program of $\mathit{nestedCombs}$. While the D&C version is more easy to be parallelized, it complicates element ordering. Since ordering is critical in our applications—where combinations represent hyperplanes that encode essential predictive information—we adopt the sequential version of $\mathit{nestedCombs}$ developed by He (2025b). The pseudo-code for this sequential version is shown in Algorithm 1.
@@ 5.2.2 Generating mixed polynomial degree hypersurface splitting rules
          $C_{1}=\left(\begin{array}[]{c}n\\ G-1\end{array}\right)$, $C_{2}=\left(\begin{array}[]{c}n\\ G\end{array}\right)$
@@
SiSince axis-parallel hyperplanes, general hyperplanes, and hypersurfaces can all be represented as combinations of data items, it is natural to ask whether we can construct a single generator capable of producing all these rules. A decision tree built from such a rule list would incorporate mixed splitting rules, thereby offering much greater flexibility.
Consider a list of polynomial degrees $\mathit{Ms}=\left[M_{1},M_{2},\ldots M_{l}\right]$, which corresponds to a list of integers $\mathit{Gs}=\left[G_{1},G_{2},\ldots G_{l}\right]$ ordered in ascending fashion ($G_{1}<G_{2}\ldots<G_{l}$), with $G_{l}$ the largest. We can specify $\mathit{modt}^{\mathit{\mathit{Ms}}}:\mathbb{N}\times\mathcal{D}\to\mathit{DTree}\left(\mathcal{H}^{Ms},\mathcal{D}\right)$ (where $\mathcal{H}^{Ms}=\mathcal{H}^{M_{1}}\cup\mathcal{H}^{M_{1}}\ldots\cup\mathcal{H}^{M_{1}}$ ) as
$\mathit{modt}_{M}=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{\mathit{Ms}}\circ\mathit{embed}_{\mathit{Ms}}$
where $\mathit{gen}_{\mathit{splits}}^{\mathit{Ms}}:\mathcal{D}\to\left[\mathcal{H}^{Ms}\right]$ is the generator that produces all possible splitting rules in $\mathcal{H}^{Ms}$.
Similar to the hypersurface generator, the composed function $\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{\mathit{Ms}}$ can also be fused. A key observation about Algorithm 1 is that, it can generate not only $G$-combinations but all $G^{\prime}$-combinations for $0\leq G^{\prime}\leq G$. However, the nested combination generator $\mathit{nestedCombs}\left(K,G\right)$ uses only the $G$-combinations and ignores smaller $G^{\prime}$-combination ($G^{\prime}<G$). To accomodate this, we need to modify specification 69 to the following
@@
where $\left(\mathit{Gs},\mathit{xss}\right)=\mathit{concat}\left[!\left(G,\mathit{xss}\right)\mid G\leftarrow\mathit{Gs}\right]$ and $\mathit{Gs}=\left[G_{1},G_{2},\ldots G_{l}\right]$ is an ascending list of integers ($G_{1}<G_{2}\ldots<G_{l}$), with $G_{l}$ being the largest, corresponds to a list of polynomial degree $\mathit{Ms}=\left[M_{1},M_{2},\ldots M_{l}\right]$. Specification 70 selects several $G$-combinations, such that $G\in\mathit{Gs}$, and constructs $K$-nested combinations from them.
As usual, the naive specification 70 is inefficient. He and Little (2025a) provide an efficient solution by defining the recursive pattern of $\mathit{nestedCombs}^{\prime}\left(K,\mathit{Gs}\right)$ as
$\bigg<\mathit{setEmpty}\left(G_{l}\right)\cdot\mathit{KcombsAlg}\left(K\right)\cdot\mathit{Ffst},$ (71)
@@
$\mathrm{convol}_{\text{new}}\left(\cup,\circ,\mathit{css}_{1},\mathit{css}_{2}\right)=$ (72)
$\quad\left[\mathit{concat}\left(\mathit{zipWith}\left(\circ,\mathit{init}\left(\mathit{tail}\left(\mathinner{css}\right)\right),\mathit{init}\left(\mathit{tail}\left(\mathit{css}_{2}\right)\right)\right)\right)\mid\mathinner{css}\leftarrow\mathit{inits}\left(\mathit{css}_{1}\right)\right].$
@@
Composing $\mathit{nestedCombs}^{\prime}\left(K,\mathit{Gs}\right)$ with $\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}$, we obtain the algorithm for solving the optimal decision tree problem with mixed splitting rules (MODT)
$\mathit{modt}_{K}^{\mathit{Ms}}=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}\left(K\right)\circ\mathit{nestedCombs}^{\prime}\left(K,\mathit{Gs}\right)\circ\mathit{embed}_{\mathit{Ms}}$ (73)
@@ 5.3.1 Complexity of size-constrained trees
Furthermore, the average-case complexity of $\mathit{sodt}_{\text{rec}}$ is substantially lower than its theoretical upper bound. The worst-case complexity of $\mathit{sodt}_{\text{rec}}$ occurs only when the tree forms a “chain” structure, which, as shown in the empirical analysis in Part II, is a rare occurrence.
@@ 5.3.2 Complexity of depth-constrained trees
$\mathcal{S}_{\text{depth}}\left(d,\mathit{xs}\right)\geq\mathcal{S}_{\text{size}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)\right)\right).$ (77)
where $\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)\right)$ denote the all possible unique splitting rules generated by $\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)$, as there might have duplicate splitting rules (for categorical data, distinct data points may have same feature values).
@@
$\mathcal{S}_{\text{size}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)\right)\right)\geq\mathcal{S}_{\text{kcombs}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)\right)\right).$ (78)
@@
$\mathcal{S}_{\text{depth}}\left(d,\mathit{xs}\right)\geq\mathcal{S}_{\text{size}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)\right)\right)\geq$ (79)
$\quad\mathcal{S}_{\text{kcombs}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)\right)\right)=\left(\begin{array}[]{c}\left|\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)\right)\right|\\ d\end{array}\right)$
As discussed, for axis parallel decision tree problem, $\left|\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right|$ is the sum of all distinct features with respect to the training data list $\mathit{xs}$, which is at most $N\times D$ for a size $N$ data list in $\mathbb{R}^{D}$.
@@
In the original source code of Mazumder et al. (2022)’s Quant-BnB algorithm, there is a function count_remain_3D, which computes the combinatorial complexity of the search space, i.e., the number of remaining trees in each iteration. The initial call to count_remain_3D, calculates this complexity directly from the input data without bounding parameters. For almost all datasets we have tested, count_remain_3D computed results that is strictly smaller than $\left(\begin{array}[]{c}\left|\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)\right)\right|\\ d\end{array}\right)$.
For instance, in the bidding dataset from Mazumder et al. (2022)’s GitHub repository, we observed the numbers of unique values per dimension as $\mathit{l}=\left[452,377,3,4680,49,22,4589,72,5\right]$ for this nine-dimensional dataset. Thus, the total number of distinct splitting rules is $\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)\right)=\mathit{sum}\left(\mathit{l}\right)=10249$. A depth-3 decision tree should contain at least three splitting rules, so the search space should contain at least $\left(\begin{array}[]{c}10249\\ 3\end{array}\right)=179376727124$ possible decision trees. However, the first call to count_remain_3D (before any search-space reduction) reports “Total number of trees = 3781512036”. This contradicts our lower bound, since every $3$-combination of splitting rule should construct at least one possible decision tree with depth smaller than 3.
Our analysis suggests a potential discrepancy in the algorithmic approach used by Mazumder et al. (2022) for solving the AODT problem, or it may indicate that their work addresses a specialized variant of the ODT problem. The lack of clarity in their problem definition makes it challenging to verify their claims. Further clarification of their counting formula would be valuable to confirm the validity of their approach and advance the understanding of this problem.